# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 2.2 Finding a Relevant Direction
Let $K\geq 2,\epsilon\in(0,1)$ and $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$ for $i\in[K]$ be a multiclass linear classifier. Let $V$ be a subspace such that for all $i,j\in[K]$ with $i\neq j$ it holds that $\|(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i)}-\mathbf{w}^{(j)}\|$. Then, it holds that
@@
For simplicity, let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$ and $t_{i,j}=t_{i}-t_{j}$. We have that
@@
We can write $\mathbf{w}^{(i,j)}=(\mathbf{w}^{(i,j)})^{V}+(\mathbf{w}^{(i,j)})^{\perp V}$, with $\|(\mathbf{w}^{(i,j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i,j)}\|$ and $\|(\mathbf{w}^{(i,j)})^{\perp V}\|\geq\sqrt{1-(\epsilon^{4}/K^{8})}\|\mathbf{w}^{(i,j)}\|$. Therefore, using the independence of any two orthogonal directions of a standard Gaussian, along with A.1, we have that
@@
In this section, given Proposition 2.4, we proceed to the proof of Theorem 2.1.
@@
We show that Algorithm 1, with high probability, returns a hypothesis $h$ with 0-1 error $O(\mathrm{OPT})+\epsilon$. Let $\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)}\in\mathbb{R}^{d}$ and $t_{1}^{*},\ldots,t_{K}^{*}\in\mathbb{R}$ be the parameters of the affine functions of an optimal classifier $f^{*}$, i.e., $f^{*}(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{*(i)}\cdot\mathbf{x}+t_{i}^{*})$ and denote by $W^{*}=\mathrm{span}(\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)})$. Let $C>0$ be an absolute constant that satisfies the statement of Proposition 2.8. Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3c) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. Also let $\epsilon^{\prime}=\mathrm{poly}(\epsilon/K)$ for a sufficiently small polynomial, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
@@
For all $k_{t}=\mathrm{poly}(K/\epsilon)$, if $\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+2\epsilon$ and $N\geq d\mathrm{poly}(K,k_{t},1/\epsilon)^{k_{t}}\log(1/\delta)$, then with probability at least $1-\delta$ there exists a unit vector $\mathbf{v}^{(t)}\in V_{t+1}$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$ for some unit vector $\mathbf{w}\in W^{*}$.
@@
First, note that using a sufficiently small polynomial for $\epsilon^{\prime}$, we have that $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k_{t}})$ for all $k_{t}=\mathrm{poly}(K/\epsilon)$ for a sufficiently small constant $c$. Denote by $\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})}$ the unit vectors return by Algorithm 2 at iteration $t$, i.e., the vectors in $\mathcal{E}_{t}$. For the sake of contradiction assume that $\lvert\mathbf{u}^{(i)}\cdot\mathbf{w}\rvert\leq\epsilon^{2}/(C^{2}K^{4})$, for all unit vectors $\mathbf{w}\in W^{*}$.
Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}$ be vectors in $\mathbb{R}^{d}$. For simplicity, define $\mathbf{w}^{(i,j)}\coloneqq\mathbf{w}^{(i)}-\mathbf{w}^{(j)},\mathbf{w}^{*(i,j)}\coloneqq\mathbf{w}^{*(i)}-\mathbf{w}^{*(j)}$. By Lemma 2.14, there exists a classifier $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}^{*})$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f^{*}(\mathbf{x})]\leq\epsilon/C$, and $\mathbf{w}^{(i)}=(\mathbf{w}^{*(i)})^{\perp U},$ for all $i\in[K]$, where $U=\mathrm{span}(\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})})$.
Since $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]\leq\mathrm{OPT}+\epsilon/C,$ we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]+\epsilon$. Therefore, by applying Proposition 2.8, for $f$, with probability $1-\delta$ there exists a vector $\mathbf{u}^{(l)}\in\mathcal{E}_{t}$ such that $\mathbf{u}^{(l)}\cdot\mathbf{w}^{(i,j)}\geq\mathrm{poly}(1/K,\epsilon)\left\|(\mathbf{w}^{(i,j)})^{\perp V_{t}}\right\|$ for some $i,j\in[K]$ with $(\mathbf{w}^{(i,j)})^{\perp V_{t}}\neq\mathbf{0}$. However, since $\mathbf{w}^{(i,j)}\in U^{\perp}$ we have a contradiction, which concludes the proof. ∎
Our next claim quantifies the decrease of the potential function $\Phi_{t}$ after adding a new weight vector.
@@
Let $\beta\geq 0$. If there exists a unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and $\mathbf{w}\in W^{*}$ such that
@@
First note that $\sum_{1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\geq\beta$ because we can write $\mathbf{w}$ as $\sum_{1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})\mathbf{b}^{(i)}$. Also note that $\lvert\mathbf{b}^{(i)}\cdot\mathbf{w}\rvert\leq\left\|\mathbf{w}\right\|=1$ for all $i\in[K]$. Hence, by the Pythagorean theorem and the dual characterization of norms, we have that
$\Phi_{t+1}=\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{\perp V_{t}}\|^{2}-\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}$ $\leq\Phi_{t}-\sum_{i=1}^{K}\left\lvert\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\right\rvert^{2}$
$\leq\Phi_{t}-\left\lvert\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\right\rvert^{2}/2$
$\leq\Phi_{t}-\beta^{2}/2\;.$
@@
Since the statement of Proposition 2.8 holds for any $\epsilon^{\prime}$-partitions $\mathcal{S}_{t},t\in[T]$, independent of the choice of basis and threshold points, we can assume that all the classifiers $h_{t}$ are computed with respect to extensions of a common orthonormal basis and that all threshold points are aligned. Hence, $\mathcal{S}_{t+1}$ is a subdivision of $\mathcal{S}_{t}$. Thus, each set $S\in\mathcal{S}_{t}$ is equal to a union of sets $S_{1},\dots,S_{l}\in\mathcal{S}_{t+1}$, i.e., $S=\cup_{i=1}^{l}S_{i}$. Therefore, by the definition of $h_{t}$, we have that
$\operatorname*{\mathbf{Pr}}[h_{t+1}(\mathbf{x})\neq y,\mathbf{x}\in S]$ $=\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}[h_{t+1}(\mathbf{x})\neq y\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\sum_{i=1}^{l}\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]$
$\leq\min_{\widehat{y}\in[K]}\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}[\widehat{y}\neq y,\mathbf{x}\in\cup_{i=1}^{l}S_{i}]$
$\leq\operatorname*{\mathbf{Pr}}[h_{t}(\mathbf{x})\neq y,\mathbf{x}\in S]\;.$
Summing over all $S\in\mathcal{S}$ we conclude the statement. ∎
In the following claim, we show that using a sufficient number of samples, we can obtain a good estimate for the population piecewise constant approximation function.
Claim 2.18.
Let $\epsilon,\epsilon^{\prime},\delta\in(0,1),k,K\in\mathbb{Z}_{+}$ with $\epsilon^{\prime}\leq\epsilon/2$. Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, and let $h$ and $\widehat{h}$ be the piecewise constant approximating functions defined according to the same partition, with width $\epsilon^{\prime}$ and subspace $V$ with respect to distributions $D$ and $\widehat{D}$, respectively (Definition 2.3), where $\widehat{D}$ is the empirical distribution of $N$ i.i.d. samples from $D$. There exists $N={(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}}\log(1/\delta)$ such that with probability $1-\delta$ we have that $\operatorname*{\mathbf{Pr}}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]+\epsilon\;.$
@@
Denote by $\mathcal{S}$ the approximating partition for the functions $h,\widehat{h}$. First, note that by the definition of an approximating partition (Definition 2.2) and by A.1, we have that $\operatorname*{\mathbf{Pr}}_{D}[S]={(\epsilon^{\prime}/k)^{\Omega(k)}}$ and hence $\lvert\mathcal{S}\rvert={(k/\epsilon^{\prime})^{O(k)}}$. Therefore, by the union bound and Hoeffding’s inequality it holds that if $N\geq C\log(\lvert\mathcal{S}\rvert K/\delta)/(\epsilon/\lvert S\rvert)^{2}={(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[\mathbf{x}\in S,y=i]-\operatorname*{\mathbf{Pr}}_{D}[\mathbf{x}\in S,y=i]\rvert\leq\epsilon/(4\lvert\mathcal{S}\rvert)$ for all $S\in\mathcal{S},i\in[K]$. Hence, with probability $1-\delta$ we have that for all $S\in\mathcal{S}$ it holds $\operatorname*{\mathbf{Pr}}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]+\epsilon/(2\lvert\mathcal{S}\rvert)$. The claim follows, as the total error is simply the summation of the error over cubes and the error outside the partition which is at most $\epsilon^{\prime}$ from Definition 2.2. ∎
@@
Note that from Lines 1, 3 of Algorithm 1 we perform at most $\mathrm{poly}(K,1/\epsilon)$ iterations. Furthermore, in each iteration, we update the vector set with at most $\mathrm{poly}(K,1/\epsilon)$ vectors. Hence, it follows that $k_{t}\leq\mathrm{poly}(K,1/\epsilon)$ for all $t=1,\dots,T$.
@@
Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$ for all $t=1,\dots,T$. Using the fact that $N=d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ for a sufficiently large polynomial (Line 2 of Algorithm 1), we can apply Claim 2.15 and conclude that with probability $1-\delta$ there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t=[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$. Thus, from Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon,1/K)$. After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon,1/K)=1-T\mathrm{poly}(\epsilon,1/K)$. However, if $T$ is set to be a sufficiently large polynomial of $\epsilon$ and $K$, we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}+\epsilon$ for some $t\in\{1,\dots,T\}$. Since the error of $h_{t}$ can only be decreasing by Claim 2.17 and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]<C\cdot\mathrm{OPT}+\epsilon$.

### v2

@@ 2.2 Finding a Relevant Direction
Let $K\geq 2,\epsilon\in(0,1)$ and $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d},t_{i}\in\mathbb{R}$ for $i\in[K]$ be a multiclass linear classifier. Let $V$ be a non-empty subspace of $\mathbb{R}^{d}$ such that for all $i,j\in[K]$ with $i\neq j$ it holds that $\|(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i)}-\mathbf{w}^{(j)}\|$. Then, it holds that
@@
For simplicity, let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$ and $t_{i,j}=t_{i}-t_{j}$. Using the union bound inequality, we have that
@@
We can write $\mathbf{w}^{(i,j)}=(\mathbf{w}^{(i,j)})^{V}+(\mathbf{w}^{(i,j)})^{\perp V}$. According to the assumptions of Lemma 2.14, it holds that $\|(\mathbf{w}^{(i,j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i,j)}\|$ and $\|(\mathbf{w}^{(i,j)})^{\perp V}\|\geq\sqrt{1-(\epsilon^{4}/K^{8})}\|\mathbf{w}^{(i,j)}\|$. Therefore, using the independence of any two orthogonal directions of a standard Gaussian, along with A.1, we have that
@@
In this section, given Proposition 2.8, we proceed to the proof of Theorem 2.1. In particular, we show that Algorithm 1 returns, with probability at least $1-\delta$, a hypothesis $h$ whose 0-1 error is $O(\mathrm{OPT})+\epsilon$.
@@
Let $\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)}\in\mathbb{R}^{d}$ and $t_{1}^{*},\ldots,t_{K}^{*}\in\mathbb{R}$ be the parameters of the affine functions of an optimal classifier $f^{*}$, i.e., $f^{*}(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{*(i)}\cdot\mathbf{x}+t_{i}^{*})$ and denote by $W^{*}=\mathrm{span}(\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)})$. Let $C>0$ be the constant that satisfies the statement of Proposition 2.8. The algorithm maintains a list of vectors $L_{t}$ (updated at Line 3c), and define $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. We set $\epsilon^{\prime}=\mathrm{poly}(\epsilon,1/K)$, for a sufficiently small polynomial (Line 2) which will be quantified later, and for each $t\in[T]$ let $\mathcal{S}_{t}$ be an arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be the piecewise constant classifier defined by $h_{t}=h_{\mathcal{S}_{t}}$ as in Definition 2.3, for the distribution $D$.
To prove the correctness of Algorithm 1, we need to show that if $h_{t}$ has significant error, then the algorithm improves its current approximation $V_{t}$ to $W^{*}$. To quantify this improvement, we define the potential function
@@
Suppose that at iteration $t$ the classifier $h_{t}$ satisfies $\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+2\epsilon$, and that $\epsilon^{\prime}$ is bounded as $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k_{t}})$, for a sufficiently small absolute constant $c>0$. Then there exists sample size, $N\geq d(k_{t}/\epsilon^{\prime})^{O(k_{t})}(K/\epsilon)^{O(1)}\log(1/\delta)$, such that with probability at least $1-\delta$ there exists unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and $\mathbf{w}\in W^{*}$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$.
@@
Let $\mathcal{E}_{t+1}=\{\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})}\}$ be the set of unit vectors returned by Algorithm 2 at iteration $t+1$. For the sake of contradiction assume that for every $\mathbf{u}\in\mathcal{E}_{t+1}$ and every unit vector $\mathbf{w}\in W^{*}$, it holds that $\lvert\mathbf{u}\cdot\mathbf{w}\rvert\leq\epsilon^{2}/(C^{2}K^{4})$.
By Lemma 2.14, there exists a classifier $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}^{*})$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f^{*}(\mathbf{x})]\leq\epsilon/C$, and for every $i\in[K]$, the vector $\mathbf{w}^{(i)}$ is defined as $\mathbf{w}^{(i)}=(\mathbf{w}^{*(i)})^{\perp U},$ where $U=\mathrm{span}(\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})})$.
Since $f$ approximates $f^{*}$ up to error $\epsilon/C$, we have $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]\leq\mathrm{OPT}+\epsilon/C$. Thus, the assumption of Claim 2.15 that $\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+2\epsilon$ implies that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]+\epsilon$. Therefore, by applying Proposition 2.8 for $f$, we deduce that the returned set $\mathcal{E}_{t+1}$ should contain a vector $\mathbf{u}^{(l)}\in\mathcal{E}_{t+1}$ so that for some indices $i,j\in[K]$, $i\neq j$, it holds $\mathbf{u}^{(l)}\cdot\mathbf{w}^{(i,j)}\geq\mathrm{poly}(1/K,\epsilon)\left\|(\mathbf{w}^{(i,j)})^{\perp V_{t}}\right\|$, where $\mathbf{w}^{(i,j)}\coloneqq\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$ with $(\mathbf{w}^{(i,j)})^{\perp V_{t}}\neq\mathbf{0}$. However, since $\mathbf{w}^{(i,j)}\in U^{\perp}$ it should hold that $\mathbf{w}^{(i,j)}\cdot\mathbf{u}^{(l)}=0$, which is a contradiction. Therefore, there exists $\mathbf{u}\in\mathcal{E}_{t+1}$ and unit vector $\mathbf{w}\in W^{*}$, such that $\lvert\mathbf{u}\cdot\mathbf{w}\rvert\geq\epsilon^{2}/(C^{2}K^{4})$, which concludes the proof. ∎
Our next claim quantifies the decrease of the potential function $\Phi_{t}$ between the iterations $t$ and $t+1$.
@@
Let $\beta\geq 0$. If there exists unit vector $\mathbf{v}^{(t)}\in V_{t+1}$ and unit vector $\mathbf{w}\in W^{*}$ such that
@@
First, note that since $\{\mathbf{b}^{(i)}\}_{i=1}^{K}$ is an orthonormal basis for $W^{*}$, any vector $\mathbf{w}\in W^{*}$ can be written as $\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})\mathbf{b}^{(i)}$. Thus, we can write $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\beta$ as $\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\geq\beta$. Recall that our potential function is defined as $\Phi_{t}=\sum_{i=1}^{K}\left\|(\mathbf{b}^{(i)})^{\perp V_{t}}\right\|^{2}$. Hence, by the Pythagorean theorem, we have that
$\Phi_{t+1}=\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{\perp V_{t}}\|^{2}-\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}=\Phi_{t}-\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}\;.$
Using the dual characterization of the norm, we can further bound
$\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}\geq\sum_{i=1}^{K}\left\lvert\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\right\rvert^{2}$
Note that from Cauchy-Schwarz inequality we have that
$\sum_{i=1}^{K}\left|\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\right|^{2}\geq\left|\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\right|^{2}\;,$
where we used that $\sum_{i=1}^{K}\left|\mathbf{b}^{(i)}\cdot\mathbf{w}\right|^{2}=1$ from Pythagorean theorem. Therefore, we have
$\Phi_{t+1}$ $\leq\Phi_{t}-\left\lvert\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\right\rvert^{2}$
$\leq\Phi_{t}-\beta^{2}\;.$
@@
Since Proposition 2.8 holds for any $\epsilon^{\prime}$-partitions $\mathcal{S}_{t},t\in[T]$, independent of the choice of basis and threshold points, we can assume that all the classifiers $h_{t}$ are computed with respect to extensions of a common orthonormal basis and that all threshold points are aligned. Hence, $\mathcal{S}_{t+1}$ is a subdivision of $\mathcal{S}_{t}$. Thus, each set $S\in\mathcal{S}_{t}$ is equal to a union of sets $S_{1},\dots,S_{l}\in\mathcal{S}_{t+1}$, i.e., $S=\cup_{i=1}^{l}S_{i}$. Therefore, by the definition of $h_{t}$, we have that
$\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t+1}(\mathbf{x})\neq y,\mathbf{x}\in S]$ $=\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t+1}(\mathbf{x})\neq y\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\sum_{i=1}^{l}\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]$
$\leq\min_{\widehat{y}\in[K]}\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[\widehat{y}\neq y,\mathbf{x}\in\cup_{i=1}^{l}S_{i}]$
$\leq\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y,\mathbf{x}\in S]\;.$
Summing over all $S\in\mathcal{S}$ we conclude the proof. ∎
In the following claim, we show that if we use a sufficiently large number of samples, then with high probability we obtain an accurate estimate of the population piecewise constant approximation function.
Claim 2.18 (Concentration of the Piecewise Constant Approximation).
Let $\epsilon,\epsilon^{\prime},\delta\in(0,1)$ and $k,K\in\mathbb{Z}_{+}$ with $\epsilon^{\prime}\leq\epsilon/2$. Let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, and consider a piecewise constant approximation $h:\mathbb{R}^{d}\to[K]$ of $D$, defined with respect to an $\epsilon^{\prime}$-approximating partition of $V$ (see Definition 2.3). Let $\widehat{D}$ be the empirical distribution obtained from $N$ i.i.d. samples drawn from $D$, and let $\widehat{h}$ be a piecewise constant approximation of $\widehat{D}$ defined with respect to the same partition. Then there exists $N=(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ such that, with probability at least $1-\delta$, we have $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]+\epsilon\;.$
@@
Let $\mathcal{S}$ denote the approximating partition used to define both $h$ and $\widehat{h}$. By the definition of an approximating partition (Definition 2.2) and using A.1, we have that for any $S\in\mathcal{S}$ it holds $\operatorname*{\mathbf{Pr}}_{D}[S]=(\epsilon^{\prime}/k)^{\Omega(k)}$ which implies that $\lvert\mathcal{S}\rvert=(k/\epsilon^{\prime})^{O(k)}$. Therefore, by the union bound and Hoeffding’s inequality it holds that if $N\geq C\log(\lvert\mathcal{S}\rvert K/\delta)/(\epsilon/\lvert S\rvert)^{2}=(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability at least $1-\delta$, it holds that for every $S\in\mathcal{S}$ and every $i\in[K]$, we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[\mathbf{x}\in S,y=i]-\operatorname*{\mathbf{Pr}}_{D}[\mathbf{x}\in S,y=i]\rvert\leq\epsilon/(4\lvert\mathcal{S}\rvert)$. Hence, with probability $1-\delta$ we have that for all $S\in\mathcal{S}$ it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]+\epsilon/(2\lvert\mathcal{S}\rvert)$. The claim follows, as the total error is simply the summation of the error over cubes and the error outside the partition which is at most $\epsilon^{\prime}$ from Definition 2.2. ∎
@@
Note that from Lines 1, 3 of Algorithm 1 we perform at most ${T=\mathrm{poly}(K,1/\epsilon)}$ iterations. Moreover, in each iteration we add at most $\mathrm{poly}(K,1/\epsilon)$ vectors to our current vector set $L_{t}$, hence the dimension $k_{t}$ of $V_{t}$ satisfies $k_{t}=\mathrm{poly}(K,1/\epsilon)$ for all $t=1,\dots,T$.
@@
Assume that for every $t=1,\dots,T$ the error satisfies $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$. Note that since $k_{t}=\mathrm{poly}(K,1/\epsilon)$ for all $t\in[T]$, if we choose $\epsilon^{\prime}=\mathrm{poly}(\epsilon/K)$ for a sufficiently small polynomial, then for all $t\in[T]$ we have $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k_{t}})$ for a sufficiently small absolute constant $c>0$. Hence, using that $N=d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ for a sufficiently large polynomial (Line 2 of Algorithm 1), we can apply Claim 2.15 and conclude that with probability $1-\delta$ there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and unit vectors $\mathbf{w}^{(t)}\in W^{*}$ for $t\in[T]$ such that $\mathbf{w}^{(t)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$. By Claim 2.16, it follows that with probability at least $1-\delta$, the potential function decreases by at least $\mathrm{poly}(\epsilon,1/K)$ at each iteration, that is $\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon,1/K)$. After $T$ iterations, it obtain that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon,1/K)=1-T\mathrm{poly}(\epsilon,1/K)$. However, since $T$ is chosen to be a sufficiently large polynomial of $\epsilon$ and $K$, we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}+\epsilon$ for some $t\in[T]$. Since the error of $h_{t}$ can only decrease over the iterations (see Claim 2.17) and $h_{t}$ is close to its empirical version (Claim 2.18), we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]<C\cdot\mathrm{OPT}+\epsilon$.